# Visual Instruction Bottleneck Tuning

### v1

@@ 4.1 Setup
Model and implementation detail. We adopt LLaVA-v1.5 liu2024improved as our main baseline MLLM, where we set CLIP ViT-L/14-336px radford2021learning as a vision encoder, Vicuna-v1.5-7B vicuna2023 as an LLM, and a two-layer MLP as a projector. We follow the standard two-stage training of LLaVA liu2023visual , and replicate stage-1 for image-text alignment with the same configuration and dataset (LLaVA-pretrain-558k) of LLaVA-v1.5 liu2024improved . Then, on the LLaVA-mix-665k, we apply our Vittle objective. To validate the scalability and broad applicability, we also consider LLaVA-v1.5-13B and Prism-7B karamcheti2024prismatic . Refer to Appendix A for details and Appendix B for LLaVA-v1.5-13B and Prism-7B results, respectively.
@@ 4.2 Results
We first evaluate Vittle on object hallucination detection tasks with nine variants of POPE perturbed by visual corruptions in Figure 3. Although MLLMs trained with a standard objective and Vittle similarly suffer from perturbations, two instantiations of Vittle consistently outperform the standard objective. Interestingly, Vittle outperforms the baseline even in clean POPE (See Appendix B). We speculate that Vittle’s information control prevents the reliance on a partial feature of a single modality rahmanzadehgervi2024vision , e.g., textual feature, which is a common source of hallucination.
@@
Next, we present the validation on the open-ended QA task with 18 types of input perturbations, which are applied to visual and textual input independently or simultaneously in Figure 4. As we can see, Vittle greatly enhances performance in various perturbation datasets highlighted by green numbers that indicate the relative improvements of Vittle (F) compared to the baseline. Among the two variants of Vittle, Vittle (F) showcases better generalization under perturbations than Vittle (L), suggesting the benefits of a conservative zero-centered isotropic prior distribution to address a variety of subtle input perturbations. Next, we further explore Vittle’s robustness by evaluating varying perturbation severity. To be specific, we generate perturbations on three different degrees that determine how significantly the image or text would be changed. In Figure 5, we see that Vittle achieves better performance in general, where the margin becomes larger under severe perturbations. In summary, we observe a consistent gain by Vittle on the perturbed input setting across two tasks, which indicates that Vittle enhances the robustness to distribution shifts by pursuing the minimality of data representation.
@@
Although the main focus of Vittle is to improve the model’s robustness under distribution shifts, securing the rich multimodal understanding capability and knowledge to diverse disciplines is also crucial as an essence of MLLM. To validate the multimodal understanding with an advanced level of knowledge, we evaluate each method on four representative closed-form QA benchmark datasets covering various fields. In Table 4.2, we observe that Vittle shows competitive performance with the standard approach, which implies that Vittle can also be used as a general-purpose learning objective.
@@
Note that the regularization forced by Vittle works on the representation space to penalize the amount of information encoded in the data representations. One of the natural alternatives is to regularize the model weight directly. In Table 4.2, we compare Vittle with LoRA hu2022lora and the weight decay method (WD) as instantiations of weight-space regularization, and the results suggest that explicit regularization on weight-space does not ensure a good balance between adaptability on in-distribution and robustness to distribution shifts. The other line of alternatives is information maximization during visual instruction tuning wang2025reconstructive ; zhou2025learning , which is the exact opposite of Vittle’s design principle. We compare two recent methods of this, ROSS wang2025reconstructive and LIT zhou2025learning , with Vittle on LB-COCO and POPE under perturbations. As shown in Figure 6, although these approaches are effective in improving object hallucination detection performances, they fail to achieve competitive performance on the open-ended QA task (See Appendix B for more results). This implies the non-trivial challenge of devising a general learning objective for MLLMs that can consistently improve robustness across diverse tasks, where we can see the promise of Vittle towards broadly applicable robust instruction tuning.
@@
Figure 7 shows responses of clean queries and their visually or textually perturbed counterparts. Although the query before and after each perturbation conveys the same meaning and intention, LLaVA-v1.5 reveals volatility in its responses, whereas Vittle shows stable behavior by providing consistent responses, i.e., generating exactly the same response in the case of visual perturbation while keep focusing on the same object in a case of textual perturbation.
@@
Vittle 9.482 13.36 0.1072
Cost analysis. Vittle introduces a lightweight bottleneck layer inside of LLM that slightly increases the total number of trainable parameters (by 1.5%). One may thus wonder how Vittle’s training and inference time is compared with a bottleneck-free baseline. In Table 5, we show the wall-clock training (per iter and total), and inference time per sample. Although Vittle increases the training time up to 20% compared with baseline, its inference time is almost identical to the original model, which is a reasonable amount of cost overhead given significant gains in terms of robustness.
@@ 5 Conclusion
This work provided the first investigation on the promise of information bottleneck from the context of MLLM instruction tuning to ensure robustness of MLLM under distribution shifts. We proposed a new theoretically-grounded visual instruction tuning method, Vittle, that injects a bottleneck layer inside the LLM to induce posterior samples of internal representations that encode useful information to produce valid responses while discarding other residual information from input queries. With negligible additional cost, Vittle is easily optimized with a variational lower bound of IB and shows consistent gains in robustness in 30 types of distribution shifts while also achieving strong performance on standard benchmarks, indicating that Vittle promotes a good balance between invariance and sensitivity during representation learning.
@@
The authors thank Seongheon Park and Sean Xuefeng Du for their valuable suggestions and discussions that shaped the draft. Research is supported in part by the AFOSR Young Investigator Program under award number FA9550-23-1-0184, National Science Foundation (NSF) Award No. IIS-2237037 and IIS-2331669, Office of Naval Research under grant number N00014-23-1-2643, Schmidt Sciences Foundation, and Alfred P. Sloan Fellowship.
@@ B.1 Ablation Study
[…]
@@ 1 Failed to converge
B.2 Full Results of Pair-wise Cosine Distance
@@
B.4 Applicability to Other MLLMs
@@
Appendix C Derivation of Variational Bound for IB in MLLM
@@
where the first inequality holds given the non-negativity of $D_{\rm KL}[r(z),p(z)]$ and $p(z_{v}|x_{v},x_{t})=p(z_{v}|x_{v})$ due to causal attention in MLLM. Here, we introduce $r(z)=r(z_{v},z_{t})=r(z_{v})r(z_{t})$ as a factorizable variational approximation of the true prior for the latent representation $p(z)$.
Next, for the output-relevant term $I(Z,Y)$, we have the lower bound:
@@
where $p(x,y,z)=p(x)p(z|x)p(y|x)$ given the Markov chain assumption $Y\leftrightarrow X\leftrightarrow Z$, and $p(z|x,y)=p(z|x)$ holds given that the representation $Z$ can not directly depend on $Y$, and the entropy term of $y$, i.e., $\int-p(y)\log p(y)dy=H(Y)$, is ruled out due to its independence for optimization problem. Here, we replace the intractable $p(y|z)$ with a variational approximation $q(y|z)$ that will be parameterized by a model.
@@
Appendix D Missing Proof
D.1 Preliminary
@@
Definition D.1 (Mutual Information (MI)).
@@
Definition D.2 (Effective Mutual Information (EMI) oh2025understanding ).
@@
D.2 A New Upper Bound for Effective Mutual Information Difference
@@
Lemma D.4 (Lemma 1 from Shui et al. shui2022novel ).
@@
Lemma D.5 (Conditional entropy bound oh2025understanding ).
Let $f(x):=H(Q_{Y|x})$ and $\hat{H}(Q_{Y|x}):=\max_{x\in\mathcal{X}}H(Q_{Y|x})$, given the marginal distributions $P_{X}$ and $Q_{X}$, and conditional distributions $P_{Y|X}$ and $Q_{Y|X}$, according to Lemma D.4, we have a conditional upper bound:
@@
Similarly, given the marginal distribution $P_{X}$ and $Q_{X}$, and an MLLM $P_{\Theta}$, let $f(x):=H(P_{\Theta}(\cdot|x))$ and $\hat{H}(P_{\Theta}):=\max_{x\in\mathcal{X}}H(P_{\Theta}(\cdot|x))$, then, according to Lemma D.4, we have another conditional upper bound:
@@
Next, we should also need to formulate the relationship between JSD in the input space and JSD in the representation space, which is done through Lemma D.6.
Lemma D.6.
@@
By applying the chain rule of KLD under a deterministic map $X\rightarrow Z=f(X)$, we know that,
@@
Now we derive a new upper bound for EMID, which is defined over the representation space rather than the previous one defined over the input space oh2025understanding in Proposition D.7.
Proposition D.7 (EMID upper bound).
Let $P_{\Theta}$ be an MLLM that maps $X=\{X_{v},X_{t}\}$ to $Z=\{Z_{v},Z_{t}\}$, and then sequentially maps $Z$ to $Y_{\Theta}$. Given joint distributions $P_{XY}=P_{X}\times P_{Y|X}$ and $Q_{XY}=Q_{X}\times Q_{Y|X}$, by assuming consistent conditional distributions over $Z_{v}|Z_{t}$, $Z_{t}|Z_{v}$, and $Y|X$ between $P$ and $Q$, we have an upper bound for $\text{EMID}(P_{XY},Q_{XY};{P_{\Theta}})$ as follow,
@@
By plugging inequalities in Lemma D.5 and ineq. (18) into the ineq. (D.2) to replace the terms (A) and (B), and given the consistent conditional distribution assumption for $Y|X$, i.e., $P_{Y|X}=Q_{Y|X}$, we have a much simpler upper bound as follows,
@@
where $\hat{H}:=\max_{x\in\mathcal{X}}[H(Q_{Y|x})+H(P_{Y_{\Theta}})]$. Then, we can further replace the term $D_{\rm JS}(P_{X}||Q_{X})$ by using Lemma D.6 to get a bound defined by representation divergence as below,
@@
Meanwhile, the chain rule of KLD and the definition of JSD with our consistent conditional distributions for $Z_{v}|Z_{t}$ and $Z_{t}|Z_{v}$, one can easily show that,
@@
Appendix E Extended Literature Review
Compression for generalization.
Robustness of fine-tuned foundation models.
Appendix F Limitation and Future Work
One of the potential concerns with IB is its reliance on the quality of $Y$, i.e., a gold response to given instruction, which is usually generated by another (M)LLM. As disclosed by Yeh et al. yeh2025challenges , the existing datasets for supervised fine-tuning are quite noisy, and we cannot ensure the advantage of IB on this noisy annotation setup. Moreover, IB alone does not guarantee the generalization of samples from completely different domains and may require additional domain information du2020learning ; li2022invariant ; ZHANG2023109737 . Investigating the potential of noisy annotation setups and domain generalization setups can be interesting future research problems. Meanwhile, a well-organized representation space by IB can be helpful for representation engineering or steering methods zou2023representation ; liu2025reducing that are also worth exploring for future work.
Appendix G Impact Statement

### v2

@@ 4.1 Setup
Model and implementation detail. We adopt LLaVA-v1.5 liu2024improved as our main baseline MLLM, where we set CLIP ViT-L/14-336px radford2021learning as a vision encoder, Vicuna-v1.5-7B vicuna2023 as an LLM, and a two-layer MLP as a projector. We follow the standard two-stage training of LLaVA liu2023visual , and replicate stage-1 for image-text alignment with the same configuration and dataset (LLaVA-pretrain-558k) of LLaVA-v1.5 liu2024improved . Then, on the LLaVA-mix-665k, we apply Vittle method by inserting the posterior MLP blocks and training the whole model with our IB objective. To validate the scalability and broad applicability, we also consider LLaVA-v1.5-13B, Prism-7B karamcheti2024prismatic , LLaVA-Mini-Vicuna-7B zhang2025llavamini , and LLaVA-Llama3-8B-Instruct hanoona2024LLaVApp . Refer to Appendix A for additional details and Appendix B for the results on LLaVA-v1.5-13B and Prism-7B, respectively.
@@ 4.2 Results
We first evaluate Vittle on object hallucination detection tasks with nine variants of POPE perturbed by visual corruptions in Figure 3. Although MLLMs trained with a standard objective and Vittle similarly suffer from perturbations, two instantiations of Vittle consistently outperform the standard objective. Interestingly, Vittle outperforms the baseline even in clean POPE (See Appendix B). We speculate that Vittle’s information control prevents the reliance on a partial feature of a single modality rahmanzadehgervi2024vision , e.g., language prior long2025understanding , which is a common source of hallucination.
@@
Next, we present the validation on the open-ended QA task with 18 types of input perturbations, which are applied to visual and textual input independently or simultaneously in Figure 4. As we can see, Vittle greatly enhances performance in various perturbation datasets highlighted by green numbers that indicate the relative improvements of Vittle (F) compared to the baseline. Among the two variants of Vittle, Vittle (F) showcases better generalization under perturbations than Vittle (L), suggesting the benefits of a conservative zero-centered isotropic prior distribution to address a variety of subtle input perturbations. Next, we further explore Vittle’s robustness by evaluating varying perturbation severity. To be specific, we generate perturbations on three different degrees that determine how significantly the image or text would be changed. In Figure 5, we see that Vittle achieves better performance in general, where the margin becomes larger under severe perturbations. In summary, we observe a consistent gain by Vittle on the perturbed input setting across two tasks, which indicates that Vittle enhances the robustness to distribution shifts by pursuing the minimal sufficiency of data representation.
@@
Although the main focus of Vittle is to improve the model’s robustness under distribution shifts, securing the rich multimodal understanding capability and knowledge to diverse disciplines is also crucial as an essence of MLLM. To validate this, we evaluate each method on four representative closed-form knowledge-intensive QA benchmark datasets covering various fields. In Table 4.2, we observe that Vittle shows competitive performance with the standard approach, which implies that Vittle can also be used as a general-purpose learning objective.
@@
Note that the regularization forced by Vittle works on the representation space to penalize the amount of information encoded in the data representations. One of the natural alternatives is to regularize the model weight directly. In Table 4.2, we compare Vittle with LoRA hu2022lora and the weight decay method (WD) as instantiations of weight-space regularization, and the results suggest that explicit regularization on weight-space does not ensure a good balance between adaptability on in-distribution and robustness to distribution shifts. The other line of alternatives is information maximization during visual instruction tuning wang2025reconstructive ; zhou2025learning , which is the exact opposite of Vittle’s design principle. We compare two recent methods on this line, ROSS wang2025reconstructive and LIT zhou2025learning , with Vittle on LB-COCO and POPE under perturbations. As shown in Figure 6, although these approaches are effective in improving object hallucination detection performances, they fail to achieve competitive performance on the open-ended QA task (See Appendix B for more results). This implies the non-trivial challenge of devising a general learning objective for MLLMs that can consistently improve robustness across diverse tasks, where we can see the promise of Vittle towards broadly applicable robust instruction tuning.
@@
Figure 7 shows responses of clean queries and their visually or textually perturbed counterparts. Although the query before and after each perturbation conveys the same meaning and intention, LLaVA-v1.5 reveals volatility in its responses, whereas Vittle shows stable behavior by providing consistent responses, i.e., generating exactly the same response in the case of visual perturbation while keep focusing on the same object in the case of textual perturbation.
@@
Vittle (F) 9.482 13.36 0.1072
Cost analysis. Vittle introduces a lightweight bottleneck layer inside of LLM that slightly increases the total number of trainable parameters (by 1.5%). One may thus wonder how Vittle’s training and inference time is compared with a bottleneck-free baseline. In Table 5, we show the wall-clock training (per iter and total), and inference time per sample. Although Vittle increases the training time up to 20% compared with baseline, its inference time is almost identical to the original model, which is a reasonable amount of cost overhead given significant gains in terms of robustness. We also compare the peak memory in Table 14, showing a negligible amount of memory overhead.
Table 6: Vittle with different model backbones.
Backbone Method POPE POPE Shifts Avg.
LLaVA-Mini zhang2025llavamini Baseline 79.37 77.39
Vittle (F) 81.07 78.32
LLaVA++ hanoona2024LLaVApp Baseline 84.60 80.54
Vittle (F) 85.87 84.08
Varying MLLM specification. As Vittle is a model-agnostic learning framework without architecture-specific constraint, we now explore Vittle training on two different MLLM specifications in Table 6: (1) LLaVA-Mini that has the same vision encoder and LLM as LLaVA-v1.5 but has different architectural design by incorporating visual token compressor and pre-modality fusion layer; and (2) LLaVA-Llama3-8B-Instruct (denoted in LLaVA++) that we just replace the LLM backbone with Llama3-8B-Instruct. As we can see, Vittle consistently outperforms the standard LLaVA training in various MLLM specifications, demonstrating its generality across different model architectures (See Appendix B.5 for more results).
@@ 5 Conclusion
This work provided the first investigation on the promise of information bottleneck in the context of MLLM instruction tuning to ensure the robustness of MLLM under distribution shifts. We proposed a new theoretically-grounded visual instruction tuning method, Vittle. It injects a bottleneck layer inside the LLM to induce posterior samples of internal representations that encode useful information to produce valid responses while discarding other residual information from input queries. With negligible additional cost, Vittle is easily optimized with a variational lower bound of IB and shows consistent gains in robustness in 30 types of distribution shifts while also achieving competitive performance on standard benchmarks, indicating that Vittle promotes a good balance between invariance and sensitivity during representation learning.
Limitation and future work.
One possible concern with Vittle is its reliance on the quality of $Y$, i.e., a gold response to given instruction, which is usually generated by another LLM. As disclosed by Yeh et al. yeh2025challenges , existing datasets for supervised fine-tuning are quite noisy, and we cannot ensure the advantage of IB on this noisy annotation setup. Moreover, IB alone does not guarantee the generalization of multi-modal counterfactual samples with conflicting language prior lee2025vlind ; long2025understanding or samples from completely different domains du2020learning ; li2022invariant ; ZHANG2023109737 , and may require additional annotations. Besides, we observed that Vittle slightly hurts the optical character recognition capability of MLLM, indicating the importance of further exploration to preserve the fine-grained recognition capability while pursuing robustness to distribution shifts. Investigating the potential of noisy annotation, counterfactual/domain generalization, and fine-grained recognition setups can be interesting future research problems. Meanwhile, the noise-robust representation space achieved by IB training can be helpful for representation steering methods zou2023representation ; liu2025reducing that are also worth exploring for future work.
@@
The authors thank Seongheon Park and Sean Xuefeng Du for their valuable suggestions and discussions that shaped the draft, and thank all the NeurIPS 2025 PCs, SACs, ACs, and anonymous reviewers. In addition, Changdae Oh particularly thanks Sanghyuk Chun for his sharp feedback on the manuscript. Research is supported in part by the AFOSR Young Investigator Program under award number FA9550-23-1-0184, National Science Foundation (NSF) Award No. IIS-2237037 and IIS-2331669, Office of Naval Research under grant number N00014-23-1-2643, Open Philanthropy, Schmidt Sciences Foundation, and Alfred P. Sloan Fellowship.
@@ B.1 Ablation Study
For bottleneck target layer ablation (Figure 12 left), we swept across $\{$8, 16, 20, 24, 28, 31}\} out of 32 layers of the 7B-size LLM backbone. However, applying the bottleneck on the early layer failed to make the language modeling loss converge, so we only provided results for 24, 28, and 31 layers. We observed that intermediate layers (L24 and L28) achieve better results than the penultimate layer (L31), and L24 shows better results on POPE while L28 outperforms L24 on LB-COCO. In conclusion, applying the bottleneck to too early parts hinders shaping some shallow syntactic features that will be actively used at later parts of the layers hernandez2021low , whereas applying it to too late parts hurts output-specific alignment or formatting song2025bridging , which guide us to decide intermediate layer, i.e., 24th, as a default choice. This is in line with a recent finding that the intermediate layer of LLM matters more than the early or later layers by showing that the quality measurements of the intermediate layer representations have a stronger correlation with performance in downstream tasks skean2025layer . Although we can search different layer indices for visual and textual tokens, we leave this to future work.
@@ 1 Failed to converge
B.2 Full Results of Pair-wise Cosine Distance Comparison
@@
B.4 Further Investigation on Vittle
Peak memory allocation.
Vittle hosts MLP blocks inside the LLM backbone to model the posterior distributions of the inner representations, and compute additional learning signals, e.g., KLD between priors and posteriors. In Table 5, we reported the wall-clock runtime, and here, we compare the maximum peak memory allocation across GPU chips (8 chips during the training and 1 chip during the inference) in Table 14. We see that the memory overhead induced by Vittle is marginal across both training and inference.
Table 14: Memory comparison. We compare baseline and Vittle (F) in terms of the maximum peak memory allocation (gigabytes; GB) across GPU chips during training and inference.
Method Peak Mem GB (train) Peak Mem GB (test)
Baseline 37.55 15.62
Vittle 38.98 15.84
Table 15: Response diversity comparison. We report scores of four representative measurements of text diversity on the models’ responses to LB-COCO clean data queries.
Method Distinct-1 ($\uparrow$) Distinct-2 ($\uparrow$) CR ($\downarrow$) Hom RL ($\downarrow$)
Baseline 0.2356 0.6117 3.234 0.155
Vittle (L) 0.2413 0.6279 3.212 0.147
Vittle (F) 0.2382 0.6260 3.238 0.144
Output diversity comparison.
One may be concerned that learning a compressive representation with Vittle can hurt the diversity of the textual output, which is undesirable for a chat assistant. To investigate the output text diversity with and without bottleneck training, we evaluate with four common textual diversity metrics, Distinct n-gram li2015diversity , as well as the compression ratio (CR) and the homogeneity score driven by Rouge-L (Hom RL) shaib2024standardizing over the outputs from each model trained by the baseline method and Vittle on the LB-COCO clean dataset.
@@ B.4 Further Investigation on Vittle
B.5 Applicability to Other MLLMs
@@
Appendix C Extended Literature Review
Compression for generalization.
Robustness of fine-tuned foundation models.
Appendix D Derivation of Variational Bound for IB in MLLM
@@
where the first inequality holds given the non-negativity of $D_{\rm KL}[r(z),p(z)]$ and $p(z_{v}|x_{v},x_{t})=p(z_{v}|x_{v})$ due to causal attention in MLLM. Here, we introduce $r(z)=r(z_{v},z_{t})=r(z_{v})r(z_{t})$ as a factorizable variational approximation of the true prior for the latent representation $p(z)$. Next, for the output-relevant term $I(Z,Y)$, we have the lower bound:
@@
where $p(x,y,z)=p(x)p(z|x)p(y|x)$ given the Markov assumption $Y\leftrightarrow X\leftrightarrow Z$, and $p(z|x,y)=p(z|x)$ holds given that the representation $Z$ can not directly depend on $Y$, and the entropy term of $y$, i.e., $\int-p(y)\log p(y)dy=H(Y)$, is ruled out due to its independence for optimization problem. Here, we replace the intractable $p(y|z)$ with a variational approximation $q(y|z)$ that will be parameterized by a model. Finally, combining the lower bound of $I(Z,Y)$ and the upper bound of $I(Z,X)$ yields a variational lower bound for the IB objective as follows,
@@
Asymmetric posterior $p(z|x)$ modeling in practice.
[…]
Appendix E Missing Proof
E.1 Preliminary
@@
Definition E.1 (Mutual Information (MI)).
@@
Definition E.2 (Effective Mutual Information (EMI) oh2025understanding ).
@@
E.2 A New Upper Bound for Effective Mutual Information Difference
@@
Lemma E.4 (Lemma 1 from Shui et al. shui2022novel ).
@@
Lemma E.5 (Conditional entropy bound oh2025understanding ).
Let $f(x):=H(Q_{Y|x})$ and $\hat{H}(Q_{Y|x}):=\max_{x\in\mathcal{X}}H(Q_{Y|x})$, given the marginal distributions $P_{X}$ and $Q_{X}$, and conditional distributions $P_{Y|X}$ and $Q_{Y|X}$, according to Lemma E.4, we have a conditional upper bound:
@@
Similarly, given the marginal distribution $P_{X}$ and $Q_{X}$, and an MLLM $P_{\Theta}$, let $f(x):=H(P_{\Theta}(\cdot|x))$ and $\hat{H}(P_{\Theta}):=\max_{x\in\mathcal{X}}H(P_{\Theta}(\cdot|x))$, then, according to Lemma E.4, we have another conditional upper bound:
@@
Next, we should also need to formulate the relationship between JSD in the input space and JSD in the representation space, which is done through Lemma E.6.
Lemma E.6.
@@
By applying the chain rule of KLD under a deterministic map [^15] Note that although Vittle governs a probabilistic encoder during training, it produces deterministic output during inference by using the learned posterior mean rather than a random sample (Section 3.3). $X\rightarrow Z$, we know that,
@@
Now we derive a new upper bound for EMID, which is defined over the representation space rather than the previous one defined over the input space oh2025understanding in Proposition E.7.
Proposition E.7 (EMID upper bound).
Let $P_{\Theta}$ be an MLLM that maps $X=\{X_{v},X_{t}\}$ to $Z=\{Z_{v},Z_{t}\}$, and then subsequently maps $Z$ to $Y_{\Theta}$. Given joint distributions $P_{XY}=P_{X}\times P_{Y|X}$ and $Q_{XY}=Q_{X}\times Q_{Y|X}$, by assuming consistent conditional distributions over $Z_{v}|Z_{t}$, $Z_{t}|Z_{v}$, and $Y|X$ between $P$ and $Q$, we have an upper bound for $\text{EMID}(P_{XY},Q_{XY};{P_{\Theta}})$ as follow,
@@
By plugging inequalities in Lemma E.5 and ineq. (18) into the ineq. (E.2) to replace the terms (A) and (B), and given the consistent conditional distribution assumption for $Y|X$, i.e., $P_{Y|X}=Q_{Y|X}$, we have a much simpler upper bound as follows,
@@
where $\hat{H}:=\max_{x\in\mathcal{X}}[H(Q_{Y|x})+H(P_{Y_{\Theta}})]$. Then, we can further replace the term $D_{\rm JS}(P_{X}||Q_{X})$ by using Lemma E.6 to get a bound defined by representation divergence as below,
@@
Meanwhile, the chain rule of KLD and the definition of JSD with our consistency assumption for conditional distributions for $Z_{v}|Z_{t}$ and $Z_{t}|Z_{v}$, one can easily show below.
@@
Appendix F Impact Statement